OpenStudy (idealist10):

Find a particular solution of y"+3y'+y=(2-6x)cosx-9sinx.

OpenStudy (idealist10):

@zepdrix @Preetha @pooja195 @TheSmartOne

OpenStudy (idealist10):

@thomaster @Luigi0210 @johnweldon1993

zepdrix (zepdrix):

Hmm I forget... Our right side contains both polynomial stuff AND cosine sine. So I think our particular solution will be of the form: product of a first degree polynomial and cosines and sines,\[\rm y_p=(A+Bx)(C \cos x+D \sin x)\quad=\quad AC \cos x+AD \sin x+BC x \cos x+BD x \sin x\]Relabeling the constants for convenience,\[\large\rm y_p=A \cos x+B \sin x+ Cx \cos x+ D x \sin x\]

zepdrix (zepdrix):

And then differentiating that is going to be a paaaaain lol

zepdrix (zepdrix):

I dunno if this helps or not, but this is my work for part of the problem. It gives us a system of 4 equations and 4 unknowns which is going to be solvable. Probably easier to approach the problem using the other method... I forget what it's called, using that weird Wronskian thing 0_o But this way isn't too bad.

OpenStudy (idealist10):

Thank you for the help!